OpenStudy (anonymous):

Please help me MEDAL REWARDED! find the solution to the quadratic equation by completing the square. @ilfy214 @nincompoop @jdoe0001 @Hero @ganeshie8 @satellite73

OpenStudy (anonymous):

\[x^{2}+6x=7\] \[x^{2}+12x=-11\]

OpenStudy (anonymous):

2 equations

OpenStudy (anonymous):

yea i know

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Lets start with the first equation. x² + 6x = 7 The first step would be to find out what value you add to both sides. You take your "x" term divide it by 2 and square it. (6/2)^2 3^2 9 We would add this to both sides of the equation x² + 6x + 9 = 7 + 9 x² + 6x + 9 = 16 Can you take it from there? Or do I need to continue?

OpenStudy (anonymous):

@Dragondd320 you lost me lol take it step by step please

OpenStudy (anonymous):

the second one is similar to the first

OpenStudy (anonymous):

\[x^2+12x=-11\] think "half of 12 is 6, and six squared is 36 go right to \[(x+6)^2=-11+36=25\]

OpenStudy (anonymous):

when i did it i got -7 and 1 and idk if im corrdct

OpenStudy (anonymous):

correct*

OpenStudy (anonymous):

Okay, first state your equation. x² + 6x = 7 Now to solve by completeing the square you use a special formula: (b/2)^2

OpenStudy (anonymous):

then you get \[x+6=5\] or \[x+6=-5\] so \[x=-1\] or \[x=-11\]

OpenStudy (anonymous):

b = 6 so (6/2)^2 3^2 9

OpenStudy (anonymous):

ok @Dragondd320 i understand so far

OpenStudy (anonymous):

Now you add this to both sides. x² + 6x +9 = 7 + 6 x² + 6x + 9 = 13 The left side is a perfect square polynomial. You can factor that down like so: x² + 6x + 9 = 13 (x + 3)(x + 3) = 13 (x + 3)² = 13 Now I ask are you wanting this solved for x, or just in vertex form?

OpenStudy (anonymous):

some typo here unfortunately

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

add 9 get \[x^2+6x+9=7+9=16\] or \[(x+3)^2=16\] other not \((x+3)^2=13\) it was just a typo is all

OpenStudy (anonymous):

I'll assume you are looking for the solution to x. *You are right @satellite73* (x + 3)² = 16

OpenStudy (anonymous):

ok can we start over because im confuse

OpenStudy (anonymous):

But I continue. (x + 3)² = 16 Take the square root of both sides. (x + 3)² = 16 x + 3 = ±4 Then solve: x + 3 = ± 4 x = 3 ± 4 So x = 3 ± 4 or: x = 3+4 x = 7 x = 3-4 x = -1 x = 7 or -1

OpenStudy (anonymous):

what goes before what you just wrote because i am confused @Dragondd320

OpenStudy (anonymous):

lets solve this one quickly, then maybe it will be clear how it works \[x^2+8x=9\] half of 8 is 4 and 4 squared is 16, go right to \[(x+4)^2=9+16=25\] or just \[(x+4)^2=25\]

OpenStudy (anonymous):

taking the square root of both sides, you get \[x+4=5\iff x=1\] or \[x+4=-5\iff x=-9\]

OpenStudy (anonymous):

im confused on how you started that

OpenStudy (anonymous):

@Dragondd320

OpenStudy (anonymous):

Sorry, I got sidetracked with something. What part has you confused?

OpenStudy (anonymous):

the first and second equation that you ans @satellite73 did im confused becuase theres was different things being said can you show me how there were suppose to be written from beginning please @Dragondd320

OpenStudy (anonymous):

Step 1: State your equation x² + 6x = 7 Step 2: Complete the square (b/2)² (6/2)² 3² = 9 Add this to both sides x² + 6x + 9 = 7 + 9 x² + 6x + 9 = 16 Step 3: Factor your perfect square x² + 6x + 9 = 16 (x + 3)(x + 3) = 16 (x + 3)² = 16 Step 4: Square both sides (x + 3)² = 16 x + 3 = ±4 (the is ± because it could be positive or negative) Step 5(final): Solve for x x + 3 = ±4 x = 3 ±4 so x = 3 + 4 and x = 3 - 4 x = 3 + 4 = 7 x = 3 - 4 = -1 x = 7 and -1 How is that for clearing up purposes?

OpenStudy (anonymous):

yes and what about the second one?

OpenStudy (anonymous):

Okay follow the same steps. Step 1: State your equation x² + 12x = -11 Step 2: Complete the square (b/2)² (12/2)² 6² = 36 Add this to both sides x² + 12x + 36 = -11 + 36 x² + 12x + 36 = 25 Step 3: Factor your perfect square x² + 12x + 36 = 25 (x + 6)(x + 6) = 25 (x + 6)² = 25 Step 4: Take the square-root of both sides. (x + 6)² = 25 x + 6 = ±5 (Could be positive or negative remember that when doing square roots) Step 5: Solve for x x + 6 = ±5 x = 6 ± 5 so x = 6+5 and x = 6-5 x = 6+5 = 11 x = 6-5 = 1 x = 11 or 1 Does that all make sense?

OpenStudy (anonymous):

yes thank you sooo much

OpenStudy (anonymous):

No problem ^.^ glad to have helped